OpenStudy (anonymous):

What is 4 2/7 multiplied by 2 1/5?

OpenStudy (anonymous):

4 2/7 =30/7 2 1/5=11/5 150/35 *77/35=11550/1225 2310/245 462/49 66/7

OpenStudy (anonymous):

ok Thank you :)